# On the Trade-off between Flatness and Optimization in Distributed Learning

### v1

@@ V Simulation results
In this section, we illustrate the performance of the three algorithms on the CIFAR10 and CIFAR100 datasets across different neural network architectures. We choose the ResNet-18 [54], WideResNet-28-10 [55] and DenseNet-121 [56] as the base neural network structures for the two datasets. As for the graph structure of the multi-agent system, we use the Metropolis rule [6] to randomly generate a doubly-stochastic graph with 16 nodes, and its structure is shown in Figure 3(a). In the main text, we present the results for ResNet-18, while the results for the other two neural networks are included Appendix J. In the decentralized experiments, the full training dataset is divided into $K$ subsets, and each agent can only observe one subset. The centralized setting is the same as the traditional single-agent learning. Moreover, we simulate three different local batch sizes including $B=128,256,512$. Note that the distributed setting with local batch $B$ means the global batch is $KB$. As for the learning scheme, we rely on the linear decaying rule [54, 57], where the initial learning rate $0.2$ is divided by $10$ when the training process reaches $50\%$ and $75\%$ of the total epoch. That is, moderately small step sizes are applied first to search for flat minima, and then smaller ones are used to guarantee the stability (i.e., convergence) of algorithms.
@@
We first illustrate the flatness and optimization performance of the three algorithms on CIFAR10 and CIFAR100. On one hand, we visualize the loss landscape around the obtained models in Figure 4. To do so, we use the visualization method from [58] where for any model $w$ we compute $J(w+\alpha\boldsymbol{v})$ using some random directions $\boldsymbol{v}$ that match the norm of $w$. It can be observed from Figure 4 that the decentralized methods arrive at models that are significantly flatter than centralized models. However, the difference in terms of flatness between diffusion and consensus is subtle. In other words, diffusion and consensus converge to models with similar flatness. On the other hand, the evolution of the training loss is shown in Figure 5, from which we observe that all three methods converge after sufficient iterations. Also, consensus consistently exhibits larger training loss in the steady state than diffusion especially when the local batch $B$ is 512, while the optimization performance of diffusion and centralized are comparable. Therefore, we conclude that diffusion enables flatter models than centralized without obviously sacrificing optimization performance. Note that we also observe some loss spikes in the training loss curves of the centralized method from Figure 5. One explanation for this phenomenon is that the centralized models may be oscillating around a sharp minima that is unstable when using moderately small step sizes. This suggests intuitively that decentralized methods aid in stabilizing the training process.
@@
Figure 4: Loss visualization around the obtained models with ResNet-18
(a) B = 128
(b) B = 256
(c) B = 512
(d) B = 128
(e) B = 256
(f) B = 512
Figure 5: The evolution of training loss with ResNet-18: (a)–(c) CIFAR10; (d)–(f) CIFAR100.
We next compare the test accuracy of the three methods on CIFAR10 and CIFAR100. To better show the performance of the three methods across different settings, we also simulate the ring graph, shown in Figure 3(b), in addition to the random graph structure. Note that the ring graph can be viewed as a specific case of a random graph generated by the Metropolis rule. Thus the random graph is more general. The simulation results are shown in Table I. In each configuration, we run the simulation 3 times with different seeds, and the final results exhibit the mean of the last 10 epochs in these repetitions. We observe from Table I that diffusion consistently outperforms the other two methods for each setting. In particular, in the cases of $B=512$, consensus even performs worse than centralized on the CIFAR100 dataset which is attributed to the optimization issue demonstrated in Figure 5.
TABLE I: Test accuracy of distributed algorithms on CIFAR10 and CIFAR100 under multiple configurations with ResNet-18.
@@
TABLE II: The generalization gap between training and test accuracy with ResNet-18.
@@
The simulation results in Table I can be interpreted by the trade-off between flatness and optimization performance. Note that our simulation results do not contradict the traditional view that flatter models generalize better [31, 32, 33, 34, 35]. We show the generalization gap measuring the difference between the training and test accuracy in all experiments in Table II, which demonstrates that flatter models found by decentralized methods enable smaller generalization gap than the centralized approach. However, we emphasize that the final test accuracy is determined by both optimization and generalization performance. When $B$ is 512 with CIFAR100 dataset, even though consensus enables better generalization, it simultaneously loses too much optimization performance. Thus its final test accuracy is worse than centralized. Similarly, diffusion achieves a favorable balance between flatness and optimization, thereby exhibiting superior test accuracy than the other two methods.
@@ VI Conclusion
We analyzed the learning behavior of three popular algorithms around local minima in nonconvex environments. The results show that decentralized methods exhibit accelerated evasion from local minima in contrast to the centralized strategy. They also show that while consensus outperforms diffusion in terms of escaping ability around local minima, this feature nevertheless comes at the expense of a deteriorated optimization performance. As a result, consensus is observed to lead to lower classification accuracy than diffusion. In other words, the results in this paper highlight an important trade-off between escaping efficiency and optimization performance in the context of multi-agent learning.
Although we focused on the traditional mini-batch gradient descent implementation, one useful extension would be to consider other types of optimizers, such as stochastic GD momentum [59] and ADAM [40]. In addition, it has been observed in the single-agent case that the structure of the gradient noise can influence the escaping efficiency and stability of stochastic gradient algorithms [34, 44]. It would be useful to examine the nature of this effect in the multi-agent case.
@@ Acknowledgments
We acknowledge the assistance of ChatGPT in improving the English expressions in this paper. We also appreciate Dr. Elsa Rizk for her valuable suggestions, which helped us present the results more clearly.
Appendix A Proof for Lemma III.4: Properties associated with gradient noise
@@
$\mathds{E}[\|\boldsymbol{s}_{k,n}(\boldsymbol{w})\|^{2}|\mathcal{F}_{n-1}]=$ $\mathds{E}\left[\left\|\frac{1}{B}\sum\limits_{b}\nabla Q_{k}(\boldsymbol{w};\boldsymbol{x}_{k,n}^{b})-\nabla J_{k}(\boldsymbol{w})\right\|^{2}\Bigg{|}\mathcal{F}_{n-1}\right]$
@@
Combining (A), (A), (A) and (A), we can verify that the variance and fourth-order moment of the gradient noise is upper bounded by terms related to batch size $B$ and the difference between the current model and local minimizer denoted by $\boldsymbol{w}-w^{\star}$:
@@
Appendix B Proof for Lemma III.5
We establish the relationship between the gradient covariance and Hessian matrices. We extend the result from single-agent case [34] to multi-agent setting. For any $w\in\mathbbm{R}^{M}$, the negative log likelihood loss is expressed as
$Q_{k}(w;\boldsymbol{x}_{k})=Q_{k}(w;\boldsymbol{h}_{k},\boldsymbol{\gamma}_{k})=-\mathrm{log}\ p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w)$ (92)
where $\boldsymbol{\gamma}_{k}$ is the true label of $\boldsymbol{h}_{k}$. Then, in the stochastic risk minimization where the risk function is defined as:
$J_{k}(w)=\mathds{E}_{\boldsymbol{x}_{k}}Q_{k}(w;\boldsymbol{x}_{k})$ (93)
we have the following gradient covariance matrix:
$\bar{R}_{s}$ =$\overset{(a)}{=}\mathds{E}\left[\left(\frac{1}{K}\sum\limits_{k}(\nabla Q_{k}(w^{\star};\boldsymbol{x}_{k})-\nabla J_{k}(w^{\star}))\right)\times\left(\frac{1}{K}\sum\limits_{\ell}(\nabla Q_{\ell}(w^{\star};\boldsymbol{x}_{\ell})-\nabla J_{\ell}(w^{\star}))\right)^{\sf T}\right]$
=$\overset{(b)}{=}\frac{1}{K^{2}}\mathds{E}\left[\sum\limits_{k}\nabla Q_{k}(w^{\star};\boldsymbol{x}_{k})\nabla^{\sf T}Q_{k}(w^{\star};\boldsymbol{x}_{k})\right]$
$=\frac{1}{K^{2}}\sum\limits_{k}\mathds{E}\left[\frac{\nabla p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w)\nabla^{\sf T}p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w)}{p_{k}^{2}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w)}\right]$ (94)
where $(a)$ follows from (33), and $(b)$ follows from (79), the independence among agents and the following equality:
$\frac{1}{K}\sum\limits_{k}\nabla J_{k}(w^{\star})=\nabla J(w^{\star})=0$ (95)
As for the Hessian matrix, we have
$\bar{H}$ $=\frac{1}{K}\mathds{E}\left[\sum\limits\nabla^{2}Q_{k}(w^{\star};\boldsymbol{x}_{k})\right]$
$=-\frac{1}{K}\mathds{E}\left[\sum\limits_{k}\nabla^{2}\mathrm{log}\ p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w^{\star})\right]$
$=-\frac{1}{K}\mathds{E}\left[\sum\limits_{k}\frac{\nabla^{2}p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w^{\star})p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w^{\star})-\nabla p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w^{\star})\nabla^{\sf T}p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w^{\star})}{p_{k}^{2}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w^{\star})}\right]$
=$\overset{(a)}{=}\frac{1}{K}\sum\limits_{k}\mathds{E}\left[\frac{\nabla p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w)\nabla^{\sf T}p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w)}{p_{k}^{2}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w)}\right]$ (96)
where $(a)$ follows from the following equality:
$\frac{1}{K}\mathds{E}\sum\limits_{k}\frac{\nabla^{2}p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w^{\star})}{p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w^{\star})}$ $=\frac{1}{K}\sum\limits_{k}\mathds{E}_{\boldsymbol{h}_{k}}\mathds{E}_{\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k}}\left[\frac{\nabla^{2}p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w^{\star})}{p_{k}(\boldsymbol{\gamma}_{k}|\boldsymbol{h}_{k},w^{\star})}\right]$
[…]
$=0$ (97)
From (B) and (B), we have the exact equivalence relationship between $\bar{R}_{s}$ and $\bar{H}$ for the stochastic risk minimization:
$\bar{R}_{s}=\frac{1}{K}\bar{H}$ (98)
Strictly speaking, since the stochastic risk minimization samples the targets from all possible predictions while the empirical optimization use the specific training label for each sample, the equivalence need not hold exactly in the empirical settings. Fortunately, we can still approximate $\bar{R}_{s}$ by $\frac{1}{K}\bar{H}$ with a Monte Carlo estimate based on the training set [53, 34], i.e., $N_{k}$ is sufficiently large at all agents. Thus, in this paper, we use
$\bar{R}_{s}\approx\frac{1}{K}\bar{H}$ (99)
Appendix C Proof for Lemmas III.6 and III.7: Upper bound for the second-order error moment
The proof in this section is similar to [6, 20, 24] with 2 important differences: we now focus on nonconvex (as opposed to convex) objective functions and on the finite-horizon (as opposed to infinite-horizon) case. To examine the size of $\mathrm{ER}_{n}$, we first analyze the mean-square error associated with ${\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}$ in (42) denoted by $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$ for later use, which can be verified that is upper bounded by a term associated with $\mu$ and gradient noise.
@@
where $(a)$ follows from (35). Multiplying (C) from the left by $\mathcal{V}^{\sf T}$, the following recursion can be obtained:
@@
According to (108), the recursion (114) can be split as
@@
Note that in the centralized method, we have
@@
so that we only need to analyze the term $\bar{\boldsymbol{w}}_{n}$ for the centralized method. Here we start from decentralized methods.
Conditioning both sides of (129) on $\mathcal{F}_{n-1}$, we now analyse the second-order moments of the two terms separately. First, for $\check{\boldsymbol{w}}_{n}$, consider
@@
where $(a)$ and $(b)$ follow from Jensen’s inequality.
@@
where $(a)$ follows from (89). Combining (133), (C), (C) and (137), we obtain
@@
where $(a)$ follows from (79), $(b)$ follows from the Jensen’s inequality and (141), and $(c)$ follows from (C) and the Lipschitz condition in Assumption III.2.
Combining (139) and (C), we obtain
$\mathds{E}\|\mathcal{V}^{\sf T}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}=\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{w}}_{n}\|^{2}\\ \|\check{\boldsymbol{w}}_{n}\|^{2}\end{array}\right]\leq\left[\begin{array}[]{cc}\frac{(1+\mu L)^{2}}{1-O(\mu)}&O(\mu)\\ O(\mu^{2})&\rho(P_{\alpha})+O(\mu^{2})\end{array}\right]\left[\begin{array}[]{c}\mathds{E}\|\bar{\boldsymbol{w}}_{n-1}\|^{2}\\ \mathds{E}\|\check{\boldsymbol{w}}_{n-1}\|^{2}\end{array}\right]+\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]$ (151)

### v2

@@ V Simulation results
In this section, we compare the performance of the three distributed algorithms on CIFAR10 and CIFAR100 datasets across different neural network architectures. We consider two distinct training scenarios: (1) Random initialization. In this scenario, all models are initialized with random parameters, ensuring that they are trained from scratch. This setup allows us to assess the algorithms’ performance in a general case without any prior knowledge. (2) Pretrained initialization. In the second scenario, the decentralized methods are initialized with models pretrained by the centralized algorithm. This approach investigates how decentralized algorithms perform when starting near a local minimum obtained through centralized training, emphasizing their ability to converge to a flatter minimum. We provide a brief overview of the experimental setup here, with detailed description available in Appendix I. We choose the ResNet-18 [55], WideResNet-28-10 [57] and DenseNet-121 [58] as the base neural network structures for the two datasets. As for the graph structure of the multi-agent system, we use the Metropolis rule [6] to randomly generate a doubly-stochastic graph with 16 nodes, and its structure is shown in Figure 3(a), and the ring topology shown in Figure 3(b). In the main text, we present the results for ResNet-18, while the results for the other two neural networks are included in Appendix I. In the decentralized experiments, the full training dataset is divided into $K$ subsets, and each agent can only observe one subset. The centralized setting is the same as the traditional single-agent learning. Moreover, we simulate three different local batch sizes including $B=128,256,512$. Note that the distributed setting with local batch $B$ means the global batch is $KB$. As for the learning scheme, we rely on the piecewise learning-rate schedule [55, 56], where the initial learning rate $0.2$ is divided by $10$ when the training process reaches $50\%$ and $75\%$ of the total epoch. That is, moderately small step sizes are applied first to search for flat minima, and then smaller ones are used to guarantee the stability (i.e., convergence) of algorithms.
@@
V-A Random initialization: training from scratch.
In this section, we show the results of the first training scenario–random initialization–where all models are trained from scratch. We first illustrate the flatness and optimization performance of the three algorithms on CIFAR10 and CIFAR100. On one hand, we visualize the risk landscape around the obtained models in Figure 4. To do so, we use the visualization method from [59] where for any model $w$ we compute the risk value $J(w+\alpha\boldsymbol{v})$ using some random directions $\boldsymbol{v}$ that match the norm of $w$. It can be observed from Figure 4 that the decentralized methods arrive at models that are significantly flatter than centralized models. However, the difference in terms of flatness between diffusion and consensus is subtle. In other words, diffusion and consensus converge to models with similar flatness. On the other hand, the evolution of the training risk is shown in Figure 5, from which we observe that all three methods converge after sufficient iterations. Also, consensus exhibits larger training risk in the steady state than diffusion especially when the local batch $B$ is $512$, while the optimization performance of diffusion and centralized are comparable. Therefore, diffusion enables flatter models than centralized without obviously sacrificing optimization performance. Note that we also observe some loss spikes in the training loss curves of the centralized method from Figure 5. One explanation for this phenomenon is that the centralized models may be oscillating around a sharp minimizer that is unstable when using moderately small step sizes. This suggests intuitively that decentralized methods aid in stabilizing the training process.
@@
Figure 4: Flatness visualization of models trained from scratch with ResNet-18. We use the visualization method from [59], according to which we compute the average training risk value $J(w+\alpha\boldsymbol{v})$ over different random directions $\boldsymbol{v}$ that match the norm of $w$. Wider valleys correspond to flatter minima.
(a) CIFAR10: B = 128
(b) CIFAR10: B = 256
(c) CIFAR10: B = 512
(d) CIFAR100: B = 128
(e) CIFAR100: B = 256
(f) CIFAR100: B = 512
Figure 5: The evolution of training risk with ResNet-18 when training from scratch: (a)–(c) CIFAR10; (d)–(f) CIFAR100.
We next compare the test accuracy of the three methods on CIFAR10 and CIFAR100. To better show the performance of the three methods across different settings, we also simulate the ring graph, shown in Figure 3(b), in addition to the random graph structure. Note that the ring graph can be viewed as a specific case of a random graph generated by the Metropolis rule. Thus the random graph is more general. The simulation results are shown in Table I. In each configuration, we run the simulation 3 times with different seeds, and the final results exhibit the mean of the last 10 epochs in these repetitions. We observe from Table I that diffusion outperforms the other two methods for each configuration when training from scratch. In particular, in the cases of $B=512$, consensus even performs worse than centralized on the CIFAR100 dataset which is attributed to the optimization issue demonstrated in Figure 5.
TABLE I: Test accuracy of distributed algorithms trained from scratch on CIFAR10 and CIFAR100 under multiple configurations with ResNet-18.
@@
V-B Pretrained initialization.
In this section, we present the results of the second scenario, where consensus and diffusion methods are initialized with models pretrained by the centralized algorithm. The flatness of the models where the three methods converge to is visualized in Figure 6. Similar to the results from the first scenario shown in Figure 4, we observe from Figure 6 that decentralized methods consistently converge to flatter minima than their centralized counterpart. Notably, in this scenario, consensus converges to flatter models than diffusion.
We then illustrate the optimization performance of consensus and diffusion in Figure 7. From this, we observe that although consensus exhibits slightly higher risk values than diffusion, their difference is more subtle compared to the results of the first scenario shown in Figure 5. This indicates that the optimization issue in the consensus method is alleviated in this scenario.
Finally, we show the test accuracy of the three methods in Table II, where the final accuracy is averaged over three repetitions with different random seeds. In this scenario, models obtained by decentralized methods consistently achieve higher accuracy than those trained using the centralized algorithm. Moreover, consensus provides slightly better test accuracy than diffusion, mainly because the pretrained initialization alleviates the effect of the generally worse optimization performance by consensus.
(a) CIFAR10: B = 128
(b) CIFAR10: B = 256
(c) CIFAR10: B = 512
(d) CIFAR100: B = 128
(e) CIFAR100: B = 256
(f) CIFAR100: B = 512
Figure 6: Flatness visualization of models with ResNet-18 using pretrained initializations in decentralized methods. The visulization method is the same with Figure 4.
(a) CIFAR10: B = 128
(b) CIFAR10: B = 256
(c) CIFAR10: B = 512
(d) CIFAR100: B = 128
(e) CIFAR100: B = 256
(f) CIFAR100: B = 512
Figure 7: The evolution of training risk with ResNet-18 using pretrained initializations in decentralized methods: (a)–(c) CIFAR10; (d)–(f) CIFAR100.
TABLE II: Test accuracy of distributed algorithms with pretrained initializations on CIFAR10 and CIFAR100 under various configurations using ResNet-18.
Dataset Method Graph $128\times 16$ $256\times 16$ $512\times 16$
Consensus Random $\textbf{92.21}\pm 0.36\%$ $\textbf{91.55}\pm 0.04\%$ $\textbf{90.57}\pm 0.24\%$
Ring $\textbf{92.31}\pm 0.37\%$ $\textbf{91.64}\pm 0.13\%$ $\textbf{90.63}\pm 0.21\%$
Diffusion Random $91.79\pm 0.30\%$ $91.21\pm 0.22\%$ $90.38\pm 0.15\%$
Ring $92.03\pm 0.36\%$ $91.26\pm 0.14\%$ $90.34\pm 0.14\%$
Consensus Random $\textbf{72.08}\pm 0.31\%$ $\textbf{71.19}\pm 0.27$% $\textbf{70.69}\pm 0.33\%$
Ring $\textbf{72.23}\pm 0.30\%$ $\textbf{71.38}\pm 0.43\%$ $\textbf{70.71}\pm 0.29\%$
Diffusion Random $71.35\pm 0.46\%$ $70.24\pm 0.28\%$ $69.85\pm 0.18\%$
Ring $71.47\pm 0.33\%$ $70.43\pm 0.53\%$ $69.76\pm 0.23\%$
V-C Discussion
The simulation results in Table I and II can be interpreted by the trade-off between generalization and optimization performance. Inspired by the findings from [30, 31, 32, 33, 34] that flatter models tend to generalize better, we present the generalization gap–which measures the difference between the training and test accuracy–in Table III and IV. These results demonstrate that flatter models found by decentralized methods result in smaller generalization gap than the centralized approach.
However, we emphasize that the final test accuracy is determined by both generalization and optimization performance. In the scenario of training from scratch, when $B$ is 512 with CIFAR100 dataset, although consensus achieves better generalization, it simultaneously sacrifices too much optimization performance. Thus, its final test accuracy is worse than the centralized algorithm. Apart from this specific case ($B=512$ with CIFAR100 dataset), consensus generally achieves a favorable balance than the centralized method, resulting in higher test accuracy. Similarly, diffusion achieves a more favorable balance between flatness and optimization in the first scenario, thereby exhibiting slightly better test accuracy than the other two methods.
In the second scenario where decentralized methods are initialized at models pretrained by the centralized algorithm, consensus converges to flatter minima than the other two methods. Additionally, its optimization issue, which previously hindered its test accuracy, is alleviated in this scenario due to pretraining. As a result, consensus provides better accuracy than the other two methods.
In summary, decentralized algorithms in general deliver enhanced classification accuracy as they strike a more favorable balance between optimization and generalization compared to the centralized solution. Furthermore, a deeper comparison between consensus and diffusion reveals distinct strengths. In general, consensus demonstrates an advantage in generalization performance, whereas diffusion excels in optimization performance. These methods exhibit different trade-off abilities between optimization and flatness across different training scenarios.
TABLE III: The generalization gap between training and test accuracy with ResNet-18 in the scenario of training from scratch.
@@ V-C Discussion
TABLE IV: The generalization gap between training and test accuracy with ResNet-18 in the scenario of pretrained initialization.
Dataset Method Graph $128\times 16$ $256\times 16$ $512\times 16$
CIFAR10 Centralized 0.0892 0.0960 0.1053
Consensus Random 0.0776 0.0837 0.0893
Ring 0.0766 0.0827 0.0879
Diffusion Random 0.0819 0.0874 0.0942
Ring 0.0795 0.0869 0.0945
CIFAR100 Centralized 0.2976 0.3099 0.3126
Consensus Random 0.2789 0.2875 0.2910
Ring 0.2774 0.2855 0.2895
Diffusion Random 0.2862 0.2972 0.3006
Ring 0.2850 0.2953 0.3016
@@ VI Conclusion
We analyzed the learning behavior of three popular algorithms around local minima in nonconvex environments. The results show that higher escaping efficiency enables algorithms to favor flatter minima but comes at the expense of deteriorated optimization performance. In general, decentralized methods exhibit accelerated evasion from local minima in contrast to the centralized strategy, without significantly compromising optimization performance. This allows decentralized methods to achieve higher classification accuracy. Furthermore, the consensus strategy has a larger excess-risk value than the diffusion one. While this performance measure is “worse” from an optimization perspective, it nevertheless comes with a benefit. It means that consensus has a higher chance of escaping from a local basin and favoring flat minima compared with diffusion. Consequently, diffusion and consensus exhibit different behavior across different scenarios. In the typical scenario of training from random initial conditions, diffusion demonstrates better test accuracy due to the more favorable balance between flatness and optimization. However, in the scenario where the decentralized methods are initialized near local minima pretrained by the centralized algorithm, consensus provides better test accuracy than diffusion. Overall, the results in this paper highlight an important trade-off between escaping efficiency and optimization performance in the context of multi-agent learning.
Although we focused on the traditional mini-batch gradient descent implementation, one useful extension would be to consider other types of optimizers, such as stochastic gradient momentum [60] and ADAM [39]. In addition, it has been observed in the single-agent case that the structure of the gradient noise can influence the escaping efficiency and stability of stochastic gradient algorithms [33, 43]. It would be useful to examine the nature of this effect in the multi-agent case.
@@ Acknowledgments
We acknowledge the assistance of ChatGPT in improving the English presentation in this paper. We also thank Dr. Elsa Rizk for her helpful suggestions.
Appendix A Poofs for Lemma III.4: Properties associated with gradient noise
@@
${\mathds{E}[\|\boldsymbol{s}_{k,n}^{B}(\boldsymbol{w})\|^{2}|\mathcal{F}_{n-1}]}=$ $\mathds{E}\left[\left\|\frac{1}{B}\sum\limits_{b}\nabla Q_{k}(\boldsymbol{w};\boldsymbol{x}_{k,n}^{b})-\nabla J_{k}(\boldsymbol{w})\right\|^{2}\Bigg{|}\mathcal{F}_{n-1}\right]$
@@
Combining (A), (A), (A) and (A), we can verify that the variance and fourth-order moment of the gradient noise is upper bounded by terms related to batch size $B$ and the difference between the current model and the local minimizer denoted by $\boldsymbol{w}-w^{\star}$:
@@
Appendix B Proof for Lemmas III.5 and III.6: Upper bound for the second-order moment
The proof in this section is similar to [6, 44, 23] with 2 important differences: we now focus on nonconvex (as opposed to convex) objective functions and on the finite-horizon (as opposed to infinite-horizon) case. To examine the size of $\mathrm{ER}_{n}$, we first analyze the mean-square distance associated with ${\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}$ in (41) denoted by $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$ for later use, which can be verified that is upper bounded by terms associated with $\mu$ and gradient noise.
@@
where $(a)$ follows from (34). Multiplying (B) from the left by $\mathcal{V}^{\sf T}$, the following recursion can be obtained:
@@
then according to (99), the recursion (105) can be split as
@@
then we have
$\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}\overset{(a)}{=}\|\mathcal{V}^{\sf T}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}=\|\bar{\boldsymbol{w}}_{n}\|^{2}+\|\check{\boldsymbol{w}}_{n}\|^{2}$ (122)
where $(a)$ follows from (95). Note that in the centralized method, we have
@@
Thus, the distance to $w^{\star}$ for decentralized models includes the additional term, $\|\check{\boldsymbol{w}}_{n}\|^{2}$, compared to the centralized method. We now prove that this extra term measures the consensus distance, i.e., the distance between local models and their centroid. Consider the centroid of local models defined by
$\boldsymbol{w}_{c,n}\overset{\Delta}{=}\frac{1}{K}\sum_{k=1}\boldsymbol{w}_{k,n}$ (124)
and the associated error vector to $w^{\star}$:
$\tilde{\boldsymbol{w}}_{c,n}\overset{\Delta}{=}w^{\star}-\boldsymbol{w}_{c,n}=\frac{1}{K}\sum_{k=1}\tilde{\boldsymbol{w}}_{k,n}$ (125)
and we introduce their collection:
${{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}\overset{\Delta}{=}\mathbbm{1}\otimes{\boldsymbol{w}}_{c,n},\quad{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}\overset{\Delta}{=}\mathbbm{1}\otimes\tilde{\boldsymbol{w}}_{c,n}=\frac{1}{K}((\mathbbm{1}\mathbbm{1}^{\sf T})\otimes I_{M}){\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}$ (126)
The consensus distance is defined by the square distance between the local models and their centroid, i.e., $\|{{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}-{{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}\|^{2}$. We first analyze the common term of decentralized and centralized methods in (122). Specifically, we have
$\bar{\boldsymbol{w}}_{n}=\left(\frac{1}{\sqrt{K}}\mathbbm{1}^{\sf T}\otimes I_{M}\right){\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}=\frac{1}{\sqrt{K}}\sum_{k}\tilde{\boldsymbol{w}}_{k,n}=\sqrt{K}\tilde{\boldsymbol{w}}_{c,n}$ (127)
so that
$\|\bar{\boldsymbol{w}}_{n}\|^{2}=K\|\tilde{\boldsymbol{w}}_{c,n}\|^{2}=\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}\|^{2}$ (128)
which means that the common term of decentralized and centralized methods denoted by $\|\bar{\boldsymbol{w}}_{n}\|^{2}$ actually measures the distance of the network centroid to the local minimizer $w^{\star}$. We next analyze the relationship between $\|\check{\boldsymbol{w}}_{n}\|^{2}$ and the consensus distance, for which we have the following equality for the difference between local models and their centroid:
${{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}-{{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}={\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}-{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}\overset{(a)}{=}(I_{KM}-\frac{1}{K}(\mathbbm{1}\mathbbm{1}^{\sf T})\otimes I_{M}){\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\overset{(b)}{=}((V_{\alpha}V_{\alpha}^{\sf T})\otimes I_{M}){\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\overset{(c)}{=}(V_{\alpha}\otimes I_{M})\check{\boldsymbol{w}}_{n}$ (129)
where $(a)$ follows from (126), $(b)$ follows from (96), and $(c)$ follows from (105). Then applying the third equality of (96), we have
$\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}-{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}\|^{2}=\|(V_{\alpha}\otimes I_{M})\check{\boldsymbol{w}}_{n}\|^{2}=\|\check{\boldsymbol{w}}_{n}\|^{2}$ (130)
Substituting (128) and (130) into (122), we have
$\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}=\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}\|^{2}+\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}-{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{c,n}\|^{2}$ (131)
Thus, the mean-square distance of the network models to the local minimizer is composed of two components: the mean-square distance of the network centroid to the local minimizer and the consensus distance. Decentralized methods include the additional consensus distance term, which increases the overall distance from the local minimizer compared to the centralized method.
We now analyze the upper bounds of the mean-square distance $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$. Here we start from decentralized methods. Conditioning both sides of (120) on $\mathcal{F}_{n-1}$, we now analyze the second-order moments of the two terms separately. First, for $\check{\boldsymbol{w}}_{n}$, consider
@@
where $(a)$ and $(b)$ follow from Jensen’s inequality, and the last constant term $O(\mu^{2})$ arises from the network heterogeneity which guarantees $d\neq 0$, otherwise this term is 0.
@@
where $(a)$ follows from (88). Combining (133), (B), (B) and (137), we obtain
@@
where $(a)$ follows from (78), $(b)$ follows from the Jensen’s inequality and (141), and $(c)$ follows from (B) and the Lipschitz condition in Assumption III.2.
Combining (139) and (B), we obtain
$\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{w}}_{n}\|^{2}\\ \|\check{\boldsymbol{w}}_{n}\|^{2}\end{array}\right]\leq\left[\begin{array}[]{cc}\frac{(1+\mu L)^{2}}{1-O(\mu)}&O(\mu)\\ O(\mu^{2})&\rho(P_{\alpha})+O(\mu^{2})\end{array}\right]\left[\begin{array}[]{c}\mathds{E}\|\bar{\boldsymbol{w}}_{n-1}\|^{2}\\ \mathds{E}\|\check{\boldsymbol{w}}_{n-1}\|^{2}\end{array}\right]+\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]$ (151)